OpenStudy (anonymous):

Two metal spheres (one with radius x cm and the other with radius x+2 cm) will be melted together to create a new single sphere. Express the radius, R, of the new sphere as a function of x.

OpenStudy (whpalmer4):

Volume of a sphere is given by \[V = \frac{4}3\pi r^3\]So calculate the volume of both spheres. Add the volumes together. Now solve the volume equation for \(r\) in terms of \(V\), plug the combined volume into that equation as \(V\) and find the value of \(r\).

OpenStudy (anonymous):

so \[4/3\pi x^3 +4/3\pi (x+2)^3=v\] and then plug that in to solve for r?

OpenStudy (whpalmer4):

Yeah. You want to expand the left hand side first, though. You'll find that the 4/3 pi stuff drops out as a common factor.

OpenStudy (anonymous):

\[(4/3\pi (x^3+(x+2)^3))/(4/3\pi)=r^3\] \[\sqrt[3]{x^3+(x+2)^3}\] does that simplify further to 2x+2 or does it stay as it is?